OpenStudy (anonymous):

A ball is dropped from the top of a building. One second later, another ball is thrown from the top of the building with a velocity of 30m/s. Will the balls ever meet? If so, when and where?

OpenStudy (shubhamsrg):

Any attempt ?

OpenStudy (anonymous):

none yet. Lack of given?

OpenStudy (shubhamsrg):

Info is sufficient .

OpenStudy (shubhamsrg):

You just need to use newton;s 2nd eqn

OpenStudy (anonymous):

we have to use formulas.

OpenStudy (shubhamsrg):

Same thing :|

OpenStudy (anonymous):

how?

OpenStudy (cwrw238):

i assume you mean 'meet in the air'. it depends on the height of the building, the formula for the distance travelled is s = ut + 0.5gt^2 so in 1 second the ball that is just dropped travels 0(1) + 16(1)^2 = 16 ft so if the building is equal or less than 16ft high they will not meet in the air.

OpenStudy (cwrw238):

otherwise cosier the flight of the 2 balls: for first ball travelling for t seconds s = 16t^2 for the second ball: suppose they meet in the air then the second ball has travelled the same distance s and has been in the air for (t - 1) seconds) so s = 30(t - 1)+ 16(t - 1)^2 so 30(t - 1)+ 16(t - 1)^2 = 16t^2 30t - 30 + 16t^2 - 16t^2 - 32t + 16 = 0 -2t = 14 t = -7 which is impossible so they never meet

OpenStudy (cwrw238):

does that make sense amistre? is my method valid?

OpenStudy (amistre64):

im contemplating it :) id use 9.8 as gravity just to be sure since the velocity given is in metters

OpenStudy (amistre64):

the dropped ball has +1 seconds on it; while the throwm ball has +0 seconds to cover the same distance.

OpenStudy (amistre64):

g(t+1)^2/2 = gt^2/2 + vt g(t+1)^2 = gt^2 + 2vt gt^2 +2gt +g = gt^2 + 2vt 2gt +g = 2vt 2gt - 2vt = -g t(2g - 2v) = -g t = -g/(2g - 2v) t = g/(2v - 2g)

OpenStudy (amistre64):

9.8/2(30-9.8) is about .25 meters

OpenStudy (amistre64):

err, seconds lol

OpenStudy (amistre64):

if we use a gravity of 32 feet per second on this then its negative .. but then the dimensions dont match up

OpenStudy (cwrw238):

silly me - i used 30 ft/sec instead of 30 m/sec

OpenStudy (amistre64):

or even 9.8 m/sec :)

OpenStudy (cwrw238):

if you replace 32 ft/s by 9.8 m/s and repeat my work the answer comes to 1.24 seconds that first ball is in the air and the distance where they meet = 7.53 metres